A generic salt, AB, has a molar mass of 313 g/mol and a solubility of 6.60 gL at 25 ℃. What is the Ksp of this salt at 25 ℃? AB s Number sp
Answer #1

Ksp = [A] * [B]

as salt is AB only, one mole of each A and B will dissociate from one mole of AB.

Moles of AB = 6.60 g/L * 1 mole / 313 g/mol = 0.021 moles

[A] = [B] = 0.021 M

Ksp = 0.021 * 0.021 = 0.000441.

Hence answer is 0.000441.